Program

Venue:

• The Hill University Center is located on University Boulevard between 14th and 15th Street South.

• The Education Building is located between 13th and 14th Street South behind a little park on University Boulevard.

• The Marshall Center is located on 11th Street South (next to the Best Western Hotel).

• The Alys Robinson Stephens Performing Arts Center is on 10th Avenue South between 12th and 13th Street South.

• The mathematics department is on the fourth floor of Campbell Hall which is on University Boulevard

between 13th and 14th Street South. Registration: Registration will start Monday, March 25, at 7:00 PM in the Marshall Center. We will also have registration in front of the lecture hall on both Tuesday and Wednesday from 8:30 to 9:00 AM and in the Education Building on Tuesday from 1:00 to 1:30 PM. Outside these times registration will take place in the office of the mathematics department (Campbell Hall 452). The office is open from 8 AM to 5 PM. We accept checks (drawn on a US financial institution) and cash. Checks are preferred. Plenary Talks: Plenary talks will be held in the mornings at 9:00 AM and 10:30 AM in the Auditorium of the Hill University Center. Parallel Sessions: Parallel sessions are in the afternoons between 1:30 and 6:00 PM. All parallel sessions take place in the Education Building. Informal Gatherings: Each night, starting on Monday, March 25, with the exception of Wednesday, March 27, there will be informal gatherings with complimentary beer, crackers and cheese. These gatherings will take place in the Marshall Center and they will start at 7:00 PM. On Thursday night (March 28) a Southern Style BBQ will be served instead of only crackers and cheese. Reception: There will be a reception with heavy hors d’oeuvres (including wine and beer) on Wednesday, March 27, in the lobby of the Alys Robinson Stephens Performing Arts Center. During the reception, the Goodfellows will provide background jazz music. There is an extra fee of $25 for the reception. It starts at 7:00 PM.

Monday, March 25

7:00 PM

MC  Informal gathering at the Marshall Center. Registration.

Tuesday, March 26

9:00 AM

(Hill) Michael Aizenman: Localization in Schrödinger Dynamics with Random Potentials

10:30 AM

(Hill) Svetlana Jitomirskaya: Nonperturbative localization.

1:30 PM

131  Rafael del Rio Castillo: Coexistence of spectra in rank-one perturbation problems
133  Gang Bao: TBA
145  Thomas Chen: Infrared renormalization and infraparticle states in QED
146  Daphne Gilbert: Spectral analysis of a class of singular differential operators
152  Maria Hoffmann-Ostenhof: A geometrical version of Hardy’s inequality

2:00 PM 131 Richard Brown: Some recent results, applications, and open problems concerning Opial’s inequality

133  Jean-Pierre Fouque: Time-reversal mirror techniques in the regime of separation of scales
145  David Damanik: Quantum Dynamical Bounds for One-Dimensional Quasicrystals

146 Pavlo Krokhmal: An approach to solving boundary-value problems of elasticity in coordinate systems with incomplete separation of variables

152  Thomas Hoffmann-Ostenhof: Eigenvalues in Symmetry Subspaces

2:30 PM

131  Robert Sims: Fractional Moment Methods for Finite Rank Perturbations

133 Jeong-Rock Yoon: On a nonlinear PDE arising in magnetic resonance electrical impedance tomography (MREIT)

145 Rainer Hempel: On the discrete spectrum of Schrödinger operators with strong magnetic fields of compact support.

146 Igor Tralle: Exact Solution of the Dirichlet Problem for Nonlinear Poisson Equation on the Plane

152  Bernhard Kawohl: Anti-eigenvalues and a conjecture of McKenna and Walter

3:00 PM

131  Mayumi Ohmiya: Darboux-Lamé equation and isomonodromic deformation on the torus
133  Lin Ji: Stiffness Identification in Biological Tissue
145  Dirk Hundertmark: A diamagnetic inequality for semigroup differences

146 Bongsoo Ko : Multiplicity of Unstable Nontrivial Solutions in Semilinear Elliptic Indefinite Weight Boundary Value Problems

152  Semjon Vugalter: Enchanced Binding in non-relativistic QED

4:00 PM 131 Antonio Hernandez: Bifurcation of relative equilibria in symmetric Hamiltonian systems at singular momentum values

133  Lizabeth Rachele: Inverse Problems for Elastic Media
145  Horst Knorrer: Asymmetric Fermi Surfaces for Magnetic Schrödinger Operators

146 Yuri Podlipenko: Minimax approach to solving some inverse problems for Maxwell equations

152 Alexander Kozhevnikov: On a Complete Scale of Isomorphisms for Elliptic and Parabolic Pseudodifferential Boundary-Value Problems

4:30 PM

131  Pavel Kurasov: Inverse scattering problem for quantum graphs
133  Chris Stolk: Constructing artifact free image families in seismic inversion
145  Heinz Siedentop: The Energy of Electrons in Heavy Atoms According to Jansen and Hess

146 Igor Popov: Asymptotics of bound states, bands and resonances for laterally coupled quantum waveguides and layers

152  Ivan Veselic: Wegner estimate with local continuity requirements on the coupling constants

5:00 PM 131 Zdzislaw Trzaska: Chaos and the coexistence of attractors in a model system of coupled difference equations

133  no talk
145  Daniel Lenz: Cantorspectrum for one-dimensional quasicrystals
146  no talk
152  Michael Hitrik: Trace distributions and heat invariants for Schrödinger operators

5:30 PM

131  Mikhail Shchukin: Homogeneous C*-algebras
133  no talk

145 Saidakhmat Lakaev: The Efimov effect of three-particle Schrödinger operators on a lattice. Asymptotics for the number of eigenvalues

146  no talk
152  Rafael Benguria: The Lane-Emden Equation revisited

7:00 PM

MC  Informal gathering at the Marshall Center.

Wednesday, March 27

9:00 AM (Hill)

Johannes Sjoestrand: Some recent progress on spectral asymptotics for non-self-adjoint operators

10:30 AM

(Hill) Ricardo Weder: The Time-Dependent Approach to Inverse Scattering

1:30 PM

131  Oleg Marichev: The website functions.wolfram.com
133  Robert Seiringer: Proof of Bose-Einstein Condensation for Dilute Trapped Gases
145  W. Desmond Evans: On the Zero modes of Pauli and Dirac operators
146  Taalaibek Imanaliev: Method of Supplementary Argument
152  Isabelle Catto: On a Dirac-Fock type models for crystals

2:00 PM 131 Ronald Mickens: Analysis of a van der Pol Type Differential Equation Having a Nonpolynomial Elastic Term

133  Marcel Griesemer: Asymptotic Completeness for Compton Scattering
145  Evans Harrell: Gap estimates for Schrödinger operators depending on curvature
146  Abdunabi Abdukarimov: On the Cauchy problem for first-order elliptic systems

152 Thomas Ostergaard Sorensen: Structure and regularity of molecular eigenfunctions and electron densities

2:30 PM

131  Kim Tuan Vu : Paley-Wiener Theorem and Sturm-Liouville Transforms
133  Laszlo Erdos: Derivation of the nonlinear Schrödinger equation with Coulomb potential
145  Edgardo Stockmeyer: The multi particle Brown-Ravenhall atom
146  Alexander Komech: On attractors of nonlinear Hamiltonian equations
152  Soeren Fournais: The total magnetic moment of large atoms in strong magnetic fields

3:00 PM

131  Jiang Wei: On the Stabilizability of Systems with Control Delay Via Output Feedback
133  Bruno Nachtergaele: From Schrödinger dynamics to the Euler equations

145 Lotfi Hermi: On Harrell-Stubbe Type Inequalities for the Discrete Spectrum of a Self-Adjoint Operator

146  Mahmadiyor Yahshiboev: Fractional integro-differentiation by Chen-Hadamard
152  Michael Pang: An alternative proof of Pascu’s theorem on hot spots

4:00 PM

131  Ivan Avramidi: Heat Kernel Asymptotics of Zaremba Boundary Value Problem

133 Gisele Goldstein: The n-dimensional Heat Equation with Nonlinear Generalized Wentzel Boundary Conditions

145 Horst Behncke: The Spectrum of Differential Operators with Almost Constant Coefficients

146  Konstantin Makarov: On a Subspace Perturbation Problem

152 Serguei Denissov: The theory of orthogonal polynomials and spectral analysis of Schrödinger operators

4:30 PM

131  Ketill Ingolfsson: Notes on the spontaneous decay
133  Jerry Goldstein: Instantaneous blow-up for heat equations
145  Andreas M. Hinz: Delocalization for Schrödinger operators

146 Alexander K. Motovilov: Factorization Problem for Operator-Valued Herglotz Functions: Geometric Approach

152 Suzanne Riehl: Consequences of the Connection Formulae for Sturm-Liouville Spectral Functions 5:00 PM 131 Martin Juras: Equations in involution and intermediate integrals for a scalar second-order hyperbolic PDE in the plane

133 Tomio Umeda: Asymptotic properties of the generalized eigenfunctions of relativistic Schrödinger operators

145  B Malcolm Brown: Computing the spectrum of the one-dimensional p-Laplacian
146  Vadim Kostrykin: Perturbation of Spectral Projections and the Riccati Equation
152  Vadim Tkachenko: 1d periodic differential operator of order 4

5:30 PM

131  Rathinasamy Sakthivel: Existence Of Solutions Of Differential Equations
133  Luc Rey-Bellet: Fluctuations of entropy production in anharmonic chains

145 Kevser Ozden Koklu: The Asymptotic Formula For The Eigenvalues Of The Second Order Partial Differential Equation With Operator Coefficient

146  Abdumalik Rakhimov: On the localization of the spectral expansions of distributions

152 Alexei Rybkin: On a new transformation of the one-dimensional Schrödinger equation and higher order WKB-type asymptotics

7:00 PM

MC  Informal gathering at the Marshall Center.

Thursday, March 28

9:00 AM

(Hill) Ari Laptev: Negative spectrum of Schrödinger operators

10:30 AM

(Hill) Claude-Alain Pillet: Spectral Theory of Open Quantum Systems

7:00 PM

MC  Informal gathering at the Marshall Center.

Friday, March 29

9:00 AM

(Hill) Elliott Lieb: Stability of a Model of Relativistic Quantum Electrodynamics

10:30 AM

(Hill) Joel Lebowitz: Complex Time Evolution of Simple Quantum Systems

1:30 PM

131  C. Maeve McCarthy: Torsional Waves in Functionally Graded Isotropic Elastic Rods
133  Susan Friedlander: A dyadic model for non-Newtonian fluid equations
145  Michael Solomyak: Schrödinger operator on homogeneous metric trees: spectrum in gaps
146  katarzyna saxton: Nonlinear balance laws in low temperature heat propagation

152 Michael Loss: A Bound on Binding Energies and Mass Renormalization in Models of Quantum Electrodynamics

2:00 PM 131 Karl Michael Schmidt: Eigenvalue asymptotics of perturbed periodic Dirac systems in the slow-decay limit

133  Natasa Pavlovic: Finite time blow-up for a dyadic model of the Euler equations

145 Leonid Friedlander: On the spectrum of second order elliptic differential operators with periodic coefficients

146  Almut Burchard: On the Cauchy problem for a dynamical Euler’s elastica
152  oliver knill: On a spectral notion of integrability

2:30 PM 131 Kyril Tintarev: Concentration compactness and nonlinear subelliptic problems without compactness

133  Mikhail Vishik: Harmonic analysis methods for incompressible flows of an ideal fluid
145  Pavel Exner: Generalized Schrödinger operators of the graph type
146  Baljeet Singh: Wave Propagation in Cracked Thermoelastic Solid Half-Space
152  Marco Lenci: Large deviations for ideal quantum systems

3:00 PM 131 Nataliya Bantsur: A note on existence and stability of periodic and almost periodic solutions of quasilinear equations with ”maxima”

133  Fedor Nazarov: TBA

145 Vassilis Papanicolaou: Recent Developments in the Theory of the Periodic Euler-Bernoulli Equation

146  Mary Ann Horn: Modelling of Nonlinear Dynamic Linked Elastic Structures
152  Leonid Bunimovich: Walks in Rigid Environments: Continuous Limits

4:00 PM

131  Toshihiko Hoshiro: The maximal smoothing effect of dispersive equations
133  Anna L. Mazzucato: The Navier-Stokes equation in distribution spaces
145  Peter Hislop: Global Continuity of the Integrated Density of States

146 Inci Albayrak: Asymptotic expression of negative eigenvalue number of an operator differential equation with even order according to small parameter having weighted function in eigenvalue

152  Wolfgang Spitzer: A new proof of the Scott correction

4:30 PM 131 Alexandra Smirnova: Regularization of nonlinear ill-posed problems by dynamical system method

133  Yuri Latushkin: Spectrum of the linearized Euler operator
145  Rowan Killip: Trace Formulae and Tri-diagonal Matrices

146 Abdurasul Gaziev: On the singular Cauchy type integrals with a continuous density for function of many variables

152 Vitali Vougalter: Pauli operator and Aharonov-Casher theorem for measure valued magnetic fields

5:00 PM

131  Gerard Awanou: Spline Approximations of the 3D Navier-Stokes Equations
133  Radu Cascaval: Pulse Waves in Elastic Tubes

145 George Hagedorn: A Time-Dependent Born-Oppenheimer Approximation with Exponentially Small Error Estimates

146  no talk
152  Kenji Yajima: Local decay for time period Schrödinger equations

5:30 PM 131 Olesya Zhupanska: A Singular Integral Equation with Sum-Difference Kernel Arising in Contact Problems with Friction

133  no talk
145  Stephen Gustafson: Effective dynamics of magnetic vortices
146  no talk
152  Alina Suzko: Time-dependent solvable models in quantum mechanics

7:00 PM

MC  Informal gathering at the Marshall Center.

Saturday, March 30

9:00 AM

(Hill) James Glimm: Predictability and Solution Error Models for Flow in Porous Media

10:30 AM

(Hill) Tai-Ping Liu: Boltzmann equation and conservation laws in gas dynamics

1:30 PM

131  Christer Bennewitz: Paley-Wiener methods in inverse spectral theory

133 Gui-Qiang Chen: Multidimensional Transonic Shocks and Free Boundary Problems for the Euler Equations for Potential Fluids

145  Tuncay Aktosun: Wave focusing on the line
146  Julio Toloza: Exponentially Small Error Estimates of Quasiclassical Eigenvalues
152  Domokos Szasz: Local Limit Theorems and Recurrence for the Planar Lorentz Process

2:00 PM

131  Christian Remling: Inverse spectral theory for one-dimensional Schrödinger operators
133  Hailiang Liu: Concentration and Cavitation of Density in Invicid Compressible Fluid Flow
145  Michael J Gruber: Spectral theory of Schrödinger Operators with periodic magnetic field
146  Daniel Ueltschi: On the sum of eigenvalues of the discrete Laplace operator
152  Thomas Murphy: Sequences of Classical Hard-Core Collisions

2:30 PM

131  Amin Boumenir: Interpolation and Inverse Spectral Theory

133 Bo Su : Discontinuous Solutions of Hamilton-Jacobi Equations: Existence, Uniqueness and Regularity.

145 Boris Belinskiy: Stability of a general linear mechanical system with a white noise in parameters

146  Eduard Tsekanovskii: On von Neumann’s problem

152 Predrag Cvitanovic: Hopf’s last hope: Spatiotemporal chaos in terms of unstable recurrent patterns

3:00 PM

131  no talk
133  Moulay Tidriri: Analysis of coupled systems of kinetic equations and conservation laws
145  Peter Kuchment: Spectra of Neumann Laplacians in thin graph like domains

146 Rajendrasinh H. Parmar: Generalized Laguerre Polynomial Solutions for Central Potentials in Quantum Mechanics

152  no talk

4:00 PM

131  Alexander Makin: On boundary value problems for one-dimensional Schrödinger operator
133  Konstantina Trivisa: Systems of Conservation Laws in Continuum Physics

145 Carlos Villegas-Blas: The regularization of the Kepler problem and the Bargmann Transform

146  Nasiba Raupova: Stability of solving systems Differential Equations of special form in R2
152  Dimitri Gioev: Lower order terms in Szego type asymptotic formulas

4:30 PM

131  Rasul Mardiev: Invertibility of the functional operators with shift having singularity
133  David Wagner: Exotherimic Reacting Euler Equations
145  Lyonell Boulton: D-S projection methods for discrete Schrödinger operators

146 Yahyo Muhtorov: Qualitative Analysis Of Two Dimensional System Differential Equations Of Special Form

152  Alexei Iantchenko: Birkhoff normal forms for Fourier integral operators

5:00 PM

131  no talk
133  Robin Young: The p-system with Large Data
145  Christian Hainzl: One non-relativistic particle coupled to a photon field

146 Vafokul Ergashev: Qualitative Analysis of Three Dimensional System Differential Equations of Special Forms.

152  Mikhail Kovalyov: nonlinear analogues of e^{i(kx−ωt)} for integrable systems

5:30 PM

131  no talk
133  no talk
145  no talk

146 Tolibjon Buriev: Qualitative Analysis of Three Dimensional System Differential Equations of Special Forms

152  no talk

7:00 PM

MC  Informal gathering at the Marshall Center.